OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, –2) with a slope of ?

OpenStudy (anonymous):

the standard form is y - h = m(x-k) ( 3, -2) is ( h, k) so y + 2 = (3/5)(x - 3)